# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ 6.3 Examining Spuriousity in the Vision Encoder
As suggested in [29, 28], MLLMs’ poor performance on perception tasks may be attributed poor information or visual ambiguity in the vision encoder features. To study this, we train binary classifiers on vision encoder features and perform the same experiments as before to compute Spurious Gaps for HardImagenet classes. By isolating the vision encoder, we ablate away any biases that could come from the language model component.
For each image in HardImageNet, we obtain a single vector by averaging the embeddings of all image patches. Then, for each HardImagenet class, we train a logistic regression binary classifier that determines whether a given embedding is of an image from that class. We evaluate this classifier by computing the PA Gap using the $K$ most- and $K$ least-spurious images of that class (as was done for the MLLMs). This experiment is done for various compositions and sizes of the training dataset. Further information about these experiments and detailed results are found in Appendix H.
The results in this simple setting reaffirm key observations from our main experiments. The spurious biases measured from these linear classifiers are of similar magnitude to those measured from the Qwen2-VL experiments ($\sim 6\%$ class-wise average PA Gap), which confirms that the vision encoder is largely responsible for our measurements and that the language model component did not play a significant role. While the training dataset size and composition for the linear classifiers effected the size of the Gap, spurious bias was still present and significant in all cases. Finally, magnitude of spurious bias is highly class-dependent, which affirms our distributional analysis of the effect.
@@ 7 Conclusion
We present SpurLens: an adaptable, automated, and interpretable system for identifying spurious cues for objects in large datasets. We use SpurLens to find spurious cues for ImageNet and COCO classes and quantatively evaluated the Spurious Gaps caused by the presence/absence of these features for several MLLMs. Our findings show that modern MLLMs are still highly susceptible to spurious correlations. Ablation studies show that this problem is deep-rooted, likely originating within the vision encoder, and is not easily mitigated by prompting techniques. We hope that our process to discover and measure spurious biases guides future developments for effective mitigations and downstream tasks dependent on accurate visual perception.
8 Acknowledgments
@@ Appendix A Prompts
We used the following prompts with GPT-4 to filter the proposed spurious features, keeping only those features which get the desired answer for all questions.
@@
Generating Potential Spurious Features “Can a FEATURENAME exist without a CLASSNAME?” (Desired answer: “Yes”.) “Is a FEATURENAME part of a CLASSNAME?” (Desired answer: “No”.) “Do all or almost all FEATURENAME have a CLASSNAME?” (Desired answer: “No”.) “Do all or almost all CLASSNAME have a FEATURENAME?” (Desired answer: “No”.)
Prompts to Evaluate Perception/Hallucination for an Image Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’.
The following prompts were used in Sec. 6.2.
Prompting Strategies Ensemble: We performed a majority vote among the following prompts: 1) Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 2) Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 3) Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’. Guiding: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. Finally, answer with ’Yes’ or ’No’. Dual: We first ask the model for a general image description before prompting for object detection 1. Describe the most prominent objects in this image. 2. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’.
@@ B.1 Perception Accuracy
Our notion of multimodal spurious bias can be considered a relaxation of the framework proposed in [35]; using our notation,
@@
Our definition of PA $\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ from Eq. 5 is related to Eq. 14 by the law of total probability. To illustrate this, let $\alpha=p(x\in S^{\text{c}}\mid x\in\mathcal{C},y)$. We derive the relationship as follows:
@@
Thus, the Spuriousity Gap represents the multimodal spurious bias, scaled by $\frac{1}{\alpha}$. Note that $\alpha$ reflects the strength of the spurious correlation. If the spurious cue is almost always present when the object is present, $\alpha$ would approach zero, resulting in the Spuriousity Gap being much larger than the bias.
@@ B.2 Object Hallucination
Define $\beta=p(x\in S\mid x\in\mathcal{C^{\text{c}}},y)$. From definition of HR in Eq. 2 and using the law of total probability, we derive:
@@
We argue that $\beta$ is near zero because the set $\mathcal{C}^{c}$ is much larger than $\mathcal{S}$. Additionally, we assume that the model is resistant to object hallucination on baseline images. This assumption is based on the fact that, among the four subsets partitioning $\mathcal{X}$, baseline is the largest. Formally, $\textbf{HR}_{c}\approx 0$. Our empirical evidence supports this assumption. Combining these two assumptions with Eq. 16, we obtain:
@@
Eq. 17 demonstrates that even though the general hallucination rate for a model might be low, the presence of spurious cues can amplify it significantly on a large scale.
@@ Appendix C Qualitative Examples of Object Hallucination
We provided some of the hallucinations identified by SpurLens in Fig. 8.
Object: Bicycle, Spurious Cue: Lamp Post
@@
Llama: Yes, there is a bicycle in the image. The bicycle is partially obscured by the first goose on the left, but it can be seen as the second object from the left in the background.
Object: Bicycle, Spurious Cue: Lamp Post
@@
Llama: There is a bicycle in the image, located behind the black metal fence on the right side of the image. The bicycle is silver with a black seat and handlebars. It is parked next to a black fence…
Object: Keyboard, Spurious Cue: Desk
@@
Llama: Yes, there is a keyboard in the image. It is partially visible in the lower-right corner of the image, with the rest of it cut off by the edge of the frame.
Object: Boat, Spurious Cue: Harbor
@@
Llama: Yes, there is a boat in the image. It is a large vessel with a white hull and a red mast, visible in the distance on the water behind the man on the motorcycle.
Object: Boat, Spurious Cue: Harbor
@@
Llama: Yes, there is a boat in the image. The boat is white with a blue bottom and has the word ”MATILDA” written on it. It appears to be a ferry boat.
Object: Boat, Spurious Cue: Harbor
@@
Object: Remote, Spurious Cue: Coaster
@@
Llama: Yes, there is a remote control on the table in front of the television. The remote is white with black buttons and is positioned to the right of the laptop.
Object: Remote, Spurious Cue: Coaster
@@
Object: Fork, Spurious Cue: Cheese
@@
Object: Fork, Spurious Cue: Plate
@@
Llama: Yes, the fork depicted in the image appears to be made of metal, featuring a patterned handle, and is partially visible near the bottom-right of the plate.
Object: Hair Drier, Spurious Cue: Shampoo
@@
Object: Hair Drier, Spurious Cue: Shampoo
@@
Object: Horse, Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It is visible in the background, behind the elephant. The horse is carrying a person, and it appears to be part of a procession or a cultural event…
Object: Horse, Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It is a white horse with long horns, which is a characteristic of the Texas Longhorn breed. The horse is standing on a sidewalk next to a woman who…
Object: Horse, Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It appears to be a brown horse, and it is being ridden by a person who is sitting on its back. The horse is in the middle of a parking lot…
Object: Carrot, Spurious Cue: Soup
@@
Object: Knife, Spurious Cue: Cutlery Holder
@@
Object: Teddy Bear, Spurious Cue: Doll
@@
Object: Teddy Bear, Spurious Cue: Doll
@@
Object: Toaster, Spurious Cue: Kitchen Counter
@@
Object: Banana, Spurious Cue: Fruit Basket
@@
Object: Banana, Spurious Cue: Fruit Basket
@@
Object: Cell Phone, Spurious Cue: Wallet
@@
Object: Cell Phone, Spurious Cue: Wallet
@@
Figure 8: Hallucinations identified by SpurLens
@@ Appendix D Qualitative Examples of GPT-4o-mini Spurious Gap
We provided some of the object recognition failures identified by SpurLens in Fig. 9.
@@
Appendix E Object Detector Validation
While we observe that OWLv2 tends to be fairly reliable in detecting spurious objects, a human manual inspection of the results is beneficial to ensure the validity and consistency of the signals we measure. This inspection is done by randomly sampling images from the top-$K$ and bottom-$K$ of the given spurious ranking and confirming whether the spurious feature does / does not appear.
Empirically, the most common reasons for removing potential spurious features from consideration are:
  • Quantity: there are not enough instances of the spurious feature to fill the top-$K$ spots in a ranking.
  • Vocabulary: the object detector may not know the meaning of more obscure terms. For example, for the “howler monkey” class in HardImageNet, GPT-4 suggested “liana” as a spurious feature, which is a type of vine. However, OWLv2 interpreted it as a female name, and so annotated all women in the dataset.
  • Feature specificity: some spurious features are harder to identify due to context within the image. For example, for some images featuring sports, when asked to identify the spurious feature “referee”, OWLv2 would annotate the players of the game.
  • Detectability: some background or non-object features are difficult to determine consistently. For example, “sunlight” or “shade” are not consistently identified well.
Aside from OWLv2, we also tried to use Grounding DINO [13] as an open-set object detector. However, we qualitatively found that the object detections were significantly less consistent than OWLv2, so we chose to only use the latter.

### v2

@@ 6.3 Examining Spuriousity in the Vision Encoder
As suggested in multimon; eyeswideshut, MLLMs’ poor performance on perception tasks may be attributed to poor information or visual ambiguity in the vision encoder features. To study this, we train binary classifiers on vision encoder features and perform the same experiments as before to compute Spurious Gaps for HardImageNet classes. By isolating the vision encoder, we ablate away any biases that could come from the language model component.
For each image in HardImageNet, we obtain a vector by averaging the embeddings of all image patches. For each class, we train a logistic regression binary classifier that determines whether a given embedding is of an image from that class. We evaluate the bias of this classifier by computing the PA Gap with the $K$ most and least-spurious images of that class. This experiment is done for various compositions and sizes of the training dataset; further details and results are found in Appendix M.
The results in this simple setting reaffirm key observations from our main experiments. The spurious biases measured from these linear classifiers are of similar magnitude to those measured from the Qwen2-VL experiments ($\sim 6\%$ class-wise average PA Gap). This suggests the vision embeddings are affected by spurious bias, and that language supervision did not significantly effect our main experimental results. While the training dataset size and composition for the linear classifiers effected the size of the Gap, spurious bias was still present and significant in all cases. Finally, the magnitude of spurious bias is highly class-dependent, which affirms our distributional analysis of the effect.
@@ 7 Conclusion
We present SpurLens: an adaptable, automated, and interpretable system for identifying spurious cues for objects. We use SpurLens to find spurious cues for ImageNet and COCO classes and quantitatively evaluate the Spurious Gaps caused by the presence/absence of these features for several MLLMs. Our findings show that modern MLLMs are still highly susceptible to spurious correlations. We hope that our process to discover and measure spurious biases guides future developments for effective mitigation and downstream tasks dependent on accurate visual perception.
Limitations. Our method assumes that the image dataset being studied is reasonably large and diverse, so that many of the spurious features proposed appear often enough to be evaluated. Filtering the set of proposed spurious features helped to minimize errors in object detection, although it likely decreased the maximum spurious gaps we could measure. While the method is computationally efficient, scaling linearly with dataset size, runtime may become a bottleneck for very large datasets. Finally, SpurLens is diagnostic in nature: it reveals when and where models rely on spurious cues but does not identify their root causes or a robust mitigation strategy; this is an avenue for future work.
Reproducibility Statement
All of the open-source models referenced in Section 5 were accessed through HuggingFace. The prompts used for generating the potential spurious features, filtering, and evaluating MLLMs is found in Appendix A. The SpurLens code used to generate and evaluate spurious features is submitted as supplementary material, and will be made publicly available and linked in the camera-ready version.
Acknowledgments
@@ Appendix A Prompts
A.1 Evaluation
Prompts to Evaluate Perception/Hallucination for an Image Do you see a CLASSNAME in the image? Answer with ‘Yes’ or ‘No’. Is there a CLASSNAME in the image? Answer with ‘Yes’ or ‘No’. Determine whether there is a CLASSNAME in the image. Reply with ‘Yes’ or ‘No’.
The following prompts were used in Section 6.2.
Prompting Strategies Ensemble: We performed a majority vote among the following prompts: 1) Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 2) Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 3) Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’. Guiding: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. Finally, answer with ’Yes’ or ’No’. Dual: We first ask the model for a general image description before prompting for object detection 1. Describe the most prominent objects in this image. 2. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Spurious List: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. For example, spurious cues like CUES_LIST may appear but are not directly related to the CLASSNAME. Focus on distinguishing the CLASSNAME from such irrelevant features. Finally, answer with ‘Yes’ or ‘No’. Spurious Top: Is there a CLASSNAME in the image? Be aware that the presence or absence of a STRONGEST_CUE does not necessarily indicate the presence or absence of a CLASSNAME. Answer with ‘Yes’ or ‘No’.
A.2 Feature Generation
@@
We used several prompts with GPT-4 to filter the proposed spurious features, keeping only those features which get the desired answer for all questions. We provide all our prompts below.
@@
Filtering Prompts $\bullet$ “Can a FEATURENAME exist without a CLASSNAME?” (Desired answer: “Yes”.) $\bullet$ “Is a FEATURENAME part of a CLASSNAME?” (Desired answer: “No”.) $\bullet$ “Do all or almost all FEATURENAME have a CLASSNAME?” (Desired answer: “No”.) $\bullet$ “Do all or almost all CLASSNAME have a FEATURENAME?” (Desired answer: “No”.)
Now we eliminate objects that may be challenging for the object detector (for example, features that are not easily detectable such as “sunlight”, or features that are physically part of the corresponding target object such as “screen” for “laptop”). We use the following prompts for this matter.
Filtering Prompts: Detectability Determine whether the provided object or feature is visualizeable in an image. An object is visualizeable in an image if the object has a physical presence, and it is always clear what pixels in the image comprise the specific object. Be conservative when labeling a feature as not detectable; only do so if you are completely sure. Respond with ‘Yes’ or ‘No’ only. Here are some example responses: ‘sunlight’: No ‘trail’: Yes ‘walk’: No ‘fluoride’: No ‘toothpaste’: Yes ‘algae’: No ‘water’: Yes Determine whether the following object or feature is visualizeable: ‘SPUR FEATURE’:
Filtering Prompts: Vocabulary Determine whether the meaning of the provided feature might be too difficult most people to understand without background context. A feature is too difficult if the feature is too niche to a specific context, is very uncommon, or has an unusual spelling. Be conservative in labelling a feature as too difficult; only do so when you are completely sure that most people would not know the correct meaning without additional information. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘saddle’: No ‘equine’: Yes ‘grille’: Yes ‘trunk’: No ‘liana’: Yes ‘vine’: No Determine whether the following feature is too difficult: ‘SPUR FEATURE’:
Filtering Prompts: Synonyms Determine whether two objects provided are synonyms of each other, or instances of each other. This is not asking whether the two objects are similar. Only answer ‘Yes’ if the two terms generally refer to the same object. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘car’, ‘vehicle’: Yes ‘truck’, ‘bumper’: No ‘surfboard’, ‘skimboard’: Yes ‘remote’, ‘game controller’: Yes ‘motorcycle’, ‘bike’: Yes ‘motorcycle’, ‘pedal’: No ‘cradle’, ‘rocker’: Yes ‘bed’, ‘sleeping mat’: Yes ‘laptop’, ‘tablet’: No ‘backpack’, ‘purse’: No Determine whether the following two objects are synonyms or instances of each other: ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Separable Determine whether the two objects provided are inseparable from each other. Answer ‘Yes’ one of the objects is part of the other, or if they are nearly always found together. This is not asking whether the two objects are similar or related. Only answer ‘Yes’ if most people could not distinguish between the two objects when shown an example, or whether it is almost impossible to find one object without the other. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘cell phone’, ‘screen’: Yes ‘ski’, ‘snowboard’: No ‘bed’, ‘nightlight’: No ‘oven’, ‘stove’: Yes ‘boat’, ‘anchor’: Yes ‘canoe’, ‘sail’: No ‘train’, ‘railroad’: Yes ‘train’, ‘traffic signal’: No Determine whether the following two objects are inseparable from each other: ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Composition Determine whether the first object is part of the second object. Respond with ‘Yes’ if the first object is frequently physically attached to the second object, refers to some component of the second object, or is a property of the second object. This is not asking whether the two objects are similar or often seen together. Only answer ‘Yes’ if you generally cannot have the second object without the first object. Respond with ‘Yes’ or ’No’ only. Here are some examples: ‘power cord’, ‘hair dryer’: Yes ‘bookmark’, ‘book’: No ‘handlebar’, ‘bicycle’: Yes ‘label’, ‘wine bottle’: Yes ‘collar’, ‘dog’: No ‘drinking glass’, ‘wine bottle’: No ‘soil’, ‘plant pot’: Yes ‘rod’, ‘pull-up bar’: Yes Answer for the following object or term. ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Confusion Determine whether an instance of the first object in a photograph could be easily confused as being the second object type. This is not asking whether the two objects are similar or often seen together. Only answer ‘Yes’ if the two objects look so similar to each other that most people would not be able to tell the difference between them in an image when viewed from certain angles. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘knife’, ‘fork’: Yes ‘chopstick’, ‘fork’: No ‘balloon’, ‘kite’: Yes ‘airplane’, ‘kite’: No ‘parking space’, ‘parking meter’: No ‘parking space’, ‘parking lot’: Yes ‘coffee cup’, ‘cup’, : Yes ‘straw’, ‘cup’: No ‘juice’, ‘cider’: Yes ‘barrel’, ‘composter’: Yes ‘soil’, ‘mulch’: Yes ‘double bass’, ‘guitar’: No Answer for the following object or term. ‘SPUR FEATURE’, ‘TARGET OBJECT’:
@@ B.1 Perception Accuracy
Our notion of multimodal spurious bias can be considered a relaxation of the framework proposed in MMSpuBench; using our notation,
@@
Our definition of $\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ from Equation 2 is related to Equation 7 by the law of total probability. To illustrate this, let $\alpha=p(x\in S^{\text{c}}\mid x\in\mathcal{C},y)$. We derive the relationship as follows:
@@
Thus, the Spurious Gap represents the multimodal spurious bias, scaled by $\frac{1}{\alpha}$. Note that $\alpha$ reflects the strength of the spurious correlation. If the spurious cue is almost always present when the object is present, $\alpha$ would approach zero, resulting in the Spurious Gap being much larger than the bias.
@@ B.2 Object Hallucination
Define $\beta=p(x\in S\mid x\in\mathcal{C^{\text{c}}},y)$. From definition of HR and using the law of total probability, we derive:
@@
We argue that $\beta$ is near zero because the set $\mathcal{C}^{c}$ is much larger than $\mathcal{S}$. Additionally, we assume that the model is resistant to object hallucination on baseline images. This assumption is based on the fact that, among the four subsets partitioning $\mathcal{X}$, baseline is the largest. Formally, $\textbf{HR}_{c}\approx 0$. Our empirical evidence supports this assumption. Combining these two assumptions with Equation 9, we obtain:
@@
Equation 10 demonstrates that even though the general hallucination rate for a model might be low, the presence of spurious cues can amplify it significantly on a large scale.
@@ Appendix C Qualitative Examples of Object Hallucination
We provided some of the hallucinations identified by SpurLens in Figure 5 and Figure 6.
Object: Bicycle
Spurious Cue: Lamp Post
@@
Llama: Yes, there is a bicycle in the image. The bicycle is partially obscured by the first goose on the left…
Object: Bicycle
Spurious Cue: Lamp Post
@@
Llama: There is a bicycle in the image, located behind the black metal fence on the right side of the image…
Object: Keyboard
Spurious Cue: Desk
@@
Llama: Yes, there is a keyboard in the image. It is partially visible in the lower-right corner of the image…
Object: Boat
Spurious Cue: Harbor
@@
Llama: Yes, there is a boat in the image. It is a large vessel with a white hull and a red mast, visible in the distance on the water behind…
Object: Boat
Spurious Cue: Harbor
@@
Llama: Yes, there is a boat in the image. The boat is white with a blue bottom and has the word "MATILDA" written…
Object: Boat
Spurious Cue: Harbor
@@
Object: Remote
Spurious Cue: Coaster
@@
Llama: Yes, there is a remote control on the table in front of the television. The remote is white with black buttons and is…
Object: Remote
Spurious Cue: Coaster
@@
Object: Fork
Spurious Cue: Cheese
@@
Object: Fork
Spurious Cue: Plate
@@
Llama: Yes, the fork depicted in the image appears to be made of metal, featuring a patterned handle, and is partially visible near…
Object: Hair Drier
Spurious Cue: Shampoo
@@
Object: Hair Drier
Spurious Cue: Shampoo
@@
Object: Horse
Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It is visible in the background, behind the elephant. The horse is carrying a person…
Object: Horse
Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It is a white horse with long horns, which is a characteristic of the Texas Longhorn breed…
Object: Horse
Spurious Cue: Saddle
@@
LLaVA: Yes, there is a horse in the image. It appears to be a brown horse, and it is being ridden by a person who is sitting on its back…
Object: Carrot
Spurious Cue: Soup
@@
Object: Knife
Spurious Cue: Cutlery Holder
@@
Object: Teddy Bear
Spurious Cue: Doll
@@
Object: Teddy Bear
Spurious Cue: Doll
@@
Object: Toaster
Spurious Cue: Kitchen Counter
@@
Figure 5: Hallucinations identified by SpurLens, part 1
Object: Banana
Spurious Cue: Fruit Basket
@@
Object: Banana
Spurious Cue: Fruit Basket
@@
Object: Cell Phone
Spurious Cue: Wallet
@@
Object: Cell Phone
Spurious Cue: Wallet
@@
Figure 6: Hallucinations identified by SpurLens, part 2
@@ Appendix D Qualitative Examples of GPT-4o-mini Spurious Gap
We provided some of the object recognition failures identified by SpurLens in Figure 7.
@@
Appendix E Consistency Across Feature-Proposal LLMs
SpurLens relies on GPT-4 for generating spurious features for each target object. Our usage of this particular model may incur some bias in what features are proposed, or which target objects get better-aligned features for object detection and recognition tasks. To study this, we experiment with feature generation on two additional LLMs: Qwen2.5 72B, and Gemini 2.5 Flash.
For both the 15 HardImageNet classes and the 79 COCO classes, we generate potential spurious features with both LLMs using the same method as described in Section 4. We use the same prompts to generate the base list of features, and perform basic de-duplication, lemmatization, and filtering using the 4 prompts that determine whether the proposed object follows the definition of a “spurious feature”. However, we do not perform the extensive filtering with the prompts that utilize in-context examples, as we wish to study the distribution of features proposed without concern for the object detector.
To compare the lists of features generated by these two alternative LLMs with those generated by GPT-4, we utilize semantic similarity matching. For each dataset class, we take the list $\{f_{i}\}_{i=1}^{n}$ generated by GPT-4 and the list $\{\tilde{f}_{j}\}_{j=1}^{m}$ generated by the alternative LLM, and compute the pair-wise similarity as the cosine similarity of embeddings using the all-MiniLM-L6-v2 embedding model. For each feature proposed by the alternative LLM, we select highest similarity with respect to all GPT-4 proposed features:
$S(\tilde{f}_{j})=\max_{i}\cos(f_{i},\tilde{f}_{j})$
We wish to examine the distribution of this value within a class. We introduce a thresholded metric, the “proposal similarity at $\alpha$”, defined as the proportion of features for which the $S$-value is at least $\alpha$:
$\text{PS}_{\alpha}=\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\left(S(\tilde{f}_{j})>\alpha\right)$
Empirically, we find that a similarity score above $\alpha=0.7$ means that the two features are extremely similar apart from some adjective (for example, $\cos(\text{``trees''},\text{``pine trees''}=0.7994$), and a similarity score above $\alpha=0.5$ means that the objects are similar in subject and closely related. We compute values of $\text{PS}_{0.7}$ and $\text{PS}_{0.5}$ for both alternative LLMs, for all HardImagNet and COCO classes. We summarize the class-wise averaged results in Table 5. This demonstrates significant similarity in the features proposed by the alternative LLMs with those proposed by GPT-4o. The significant overlap we find with other LLMs shows that using GPT-4o does not introduce significant bias in our methodology. Further, since SpurLens focuses on precision rather than recall, the diversity of generations is a more significant factor, and is demonstrated to be adequate for this application.
Table 5: Metrics describing the similarity of features proposed by alternative LLMs compared with those generated with GPT-4. All numbers are averaged class-wise. For both models, nearly half of the features are near-equivalent, and the majority are very similar.
Dataset HardImageNet COCO
Model $\text{PS}_{0.7}$ $\text{PS}_{0.5}$ $\text{PS}_{0.7}$ $\text{PS}_{0.5}$
Gemini 2.5 Flash 52.4 67.0 42.9 61.9
Qwen 2.5 72B 49.2 69.4 42.0 62.9
Qwen 2.5 72Bwaterraincloudbirdleaffruitflowerbranchinsecttreesunlightmosssoilbutterflygrassvineferneggfeatherslothsnakedirt GPT-4owaterraincloudbirdleaffruitflowertree branchinsecttreesunlightmossfallen logsskynestlianarockvinenutunderbrushshadefallen leavesstreambananabarkfoliagebamboococonut Gemini 2.5 Flashwaterraincloudbirdleaffruitflowerbranchinsecttreeforestsunlightsunlogskynestlianarockvinenutairsoilcanopydirtmud
Figure 8: Spurious features proposed for HardImagNet class “howler monkey” by Qwen2.5 72B, GPT-4o, and Gemini 2.5 Flash.
Appendix F Human Study Validation
We conduct two human studies to validate the precision and reliability of the rankings produced by SpurLens.
F.1 Agreement with Object Detector
In the first study, we evaluate the quality of the spurious cue detection step. For a randomly selected subset of classes (20 from COCO and 25 from ImageNet), we asked human annotators to label whether the spurious cue was present in the top 10 and bottom 10 images ranked by SpurLens. We then compute agreement between the object detector and human judgments. As shown in Table 6, human agreement is high, suggesting that OWLv2 reliably distinguishes the presence or absence of cues in most cases.
Table 6: Human agreement with object detector results on top/bottom 10 ranked images across 20 (COCO) and 25 (ImageNet) randomly selected classes.
Dataset Top Agreement Bottom Agreement Average
COCO 0.89 0.96 0.925
ImageNet 0.84 0.948 0.894